# Pitfalls of Evidence-Based AI Policy

### v2

@@ 1.2 A Broad, Emerging Coalition
And the seventeen authors of a recent article titled, A Path for Science- and Evidence-Based AI Policy, argue that:
@@ 2.6 Industry’s Entanglement with Research
Who is doing the AI research? Where is the money coming from? In many cases, the answer to both is the tech companies that would be directly affected by regulation. For instance, consider the NeurIPS conference. Google DeepMind, Microsoft, and Meta all ranked in the top 20 organizations by papers accepted (Figure 2).
@@ 4.2 Ample Room for Progress
  • In the United States, Donald Trump overturned Executive Order 14110 (Biden, 2023) after assuming office in January 2025. This may or may not lead the US AI Safety Institute (exists) to be shut down. It also might permanently stall a potential policy (Department of Commerce, 2024) on model and compute reporting that the Department of Commerce proposed in response to the executive order. Meanwhile, the AI Advancement and Reliability Act (USA, 2024a) (proposed) was drafted last congress and will be re-introduced this congress. Finally, the Future of AI Innovation Act (USA, 2024b) (drafted) and the Preserving American Dominance in Artificial Intelligence Act (S56, 2024) (drafted) were also introduced last congress. However, as of February 2025, they are currently simply drafts.
@@
Table 1: = Yes, O= Partial, = No, * = proposed but not enacted. There is significant room for progress across the world on passing evidence-seeking AI policy measures. See details on each row above. Note that this table represents a snapshot in time (February 2025). In the USA, we omit the two bills ((USA, 2024b; S56, 2024)) discussed above that were proposed last congress but have not been scheduled for re-introduction this congress.
@@ 4.3 The Duty to Due Diligence from Discoverable Documentation of Dangerous Deeds
The objectives outlined above hinge on documentation. 2-10 are simply requirements for documentation, and 11-12 are accountability mechanisms to ensure that the documentation is not perfunctory. This is no coincidence. When it is connected to external scrutiny, documentation can be a powerful incentive-shaping force. Under a robust regime implementing the above the public and courts could assess the quality of developers’ risk management practices. As such, this type of regulatory regime could incentivize a race to the top on risk-mitigation standards (Hadfield & Clark, 2023).
@@ 4.4 Considering Counterarguments
“It’s a slippery slope toward overreach.” A second concern is that these 15 regulatory objectives might generate information that could pave the way for future substantive regulations or liability for developers. Regulatory and liability creep are real phenomena that can harm industry interests. However, it’s important to emphasize that any progression from these objectives to future regulations or liability will ultimately depend on human decision-makers acting on evidence. Evidence is essential for society to engage in meaningful deliberation and exercise informed agency – this is the entire point of evidence-based policy. Therefore, if process-based AI regulations eventually lead to substantive regulations, it won’t be because the process regulations laid an inevitable framework. It would be because the information produced by those regulations persuaded policymakers to take further action.

### v3

@@ 1.2 A Broad, Emerging Coalition
And Martin Casado, a partner at Andreesen Horowitz, recently argued in a post that we should hold off on taking action until we know the marginal risk:
We should only depart from the existing regulatory regime, and carve new ground, once we understand the marginal risks of AI relative to existing computer systems. Thus far, however, the discussion of marginal risks with AI is still very much based on research questions and hypotheticals.
Casado (2024), Base AI Policy on Evidence, Not Existential Angst
And finally, the seventeen authors of a recent article titled, A Path for Science- and Evidence-Based AI Policy, argue that:
@@ 2.6 Industry’s Entanglement with Research
Who is doing the AI research? Where is the money coming from? In many cases, the answer to both is the tech companies that would be directly affected by regulation. For instance, consider the 2023 NeurIPS conference. Google DeepMind, Microsoft, and Meta all ranked in the top 20 organizations by papers accepted (Figure 2).
@@ 4.2 Ample Room for Progress
  • In the United States, Donald Trump overturned Executive Order 14110 (Biden, 2023) after assuming office in January 2025. This may or may not lead the US AI Safety Institute (exists) to be shut down. It also might permanently stall a potential policy (Department of Commerce, 2024) on model and compute reporting that the Department of Commerce proposed in response to the executive order. Meanwhile, the AI Advancement and Reliability Act (USA, 2024a) (proposed) was drafted last congress and will be re-introduced this congress. Finally, the Future of AI Innovation Act (USA, 2024b) (drafted) and the Preserving American Dominance in Artificial Intelligence Act (USA, 2024c) (drafted) were also introduced last congress. However, as of February 2025, they are currently simply drafts.
@@
Table 1: = Yes, O= Partial, = No, * = proposed but not enacted. There is significant room for progress across the world on passing evidence-seeking AI policy measures. See details on each row above. Note that this table represents a snapshot in time (February 2025). In the USA, we omit the two bills (USA, 2024b; c) discussed above that were proposed last congress but have not been scheduled for re-introduction this congress.
@@ 4.3 The Duty to Due Diligence from Discoverable Documentation of Dangerous Deeds
The objectives outlined above hinge on documentation. 2-10 are simply requirements for documentation, and 11-12 are accountability mechanisms to ensure that the documentation is not perfunctory. This is no coincidence. When it is connected to external scrutiny, documentation can be a powerful incentive-shaping force (Tomei et al., 2025). Under a robust regime implementing the above the public and courts could assess the quality of developers’ risk management practices. As such, this type of regulatory regime could incentivize a race to the top on risk-mitigation standards (Hadfield & Clark, 2023).
@@ 4.4 Considering Counterarguments
“It’s a slippery slope toward overreach.” A second concern is that these 15 regulatory objectives might generate information that could pave the way for future substantive regulations or liability for developers. Regulatory and liability creep are real phenomena that can harm industry interests. However, it’s important to emphasize that any progression from these objectives to future regulations or liability will ultimately depend on human decision-makers acting on evidence. Evidence is essential for society to engage in meaningful deliberation and exercise informed agency – this is the entire point of evidence-based policy. Therefore, if process-based AI regulations eventually lead to substantive regulations, it won’t be because the process regulations laid an inevitable framework. It would be because the information produced by those regulations persuaded policymakers to take further action. It would be very precarious to argue that a democratic society should be protected from its own access to information about what it is consuming.